Calculate Karl Pearson’s Coefficient of Correlation X 10 10 11 12 12 y 5 6 4 3


1
Expert's answer
2021-01-26T03:41:04-0500

The formula for Karl's Pearson's cCoefficient is:

"r=\frac{n\sum_{i=0}^{n}x_iy_i-(\sum_{i=0}^{n}x_i)(\sum_{i=0}^{n}y_i)}{\sqrt{[n\sum_{i=0}^{n}x^2_i-(\sum_{i=0}^{n}x_i)^2][n\sum_{i=0}^{n}y^2_i-(\sum_{i=0}^{n}y_i)^2]}}"

Since there are 4 values fo y and 5 for x, we assume that the last value is for y 0.

We set n=5 and receive:

"n\sum_{i=0}^nx_iy_i=5\cdot(10\cdot5+10\cdot6+11\cdot4+12\cdot3+12\cdot0)=950"

"(\sum_{i=0}^{n}x_i)(\sum_{i=0}^{n}y_i)=(10+10+11+12+12)(5+6+4+3+0)=55\cdot18=990"

"n\sum_{i=0}^{n}x^2_i-(\sum_{i=0}^{n}x_i)^2=5(10^2+10^2+11^2+12^2+12^2)-(10+10+11+12+12)^2="

"=20"

"n\sum_{i=0}^{n}y^2_i-(\sum_{i=0}^{n}y_i)^2=5(5^2+6^2+4^2+3^2+0^2)-(5+6+4+3+0)^2=106"

"r=\frac{950-990}{\sqrt{20\cdot106}}=-0.8687"

Answer:-0.8687